Determine whether the growth (or decay) is linear or exponential, and answer the associated question.
During the worst periods of inflation in America, the price of food increased at a rate of 10% per month. If your food bill was $400 one month during this period, what was it two months later?
A) Exponential; $484.00
B) Linear; $880.00
C) Linear; $480.00
D) Exponential; $441.00

Answer by Boreal(15235)

This is exponential, because the 10% rise is added on and figured in the next calculation. It's not increasing with a fixed number but a percent, and as the number increases, the percent increases more.
it is 400*(1.1)^2=400*1.21=$484
A.